Minutes of the 9th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson

Minutes taker Mark Dennehy

Present Steven Watterson (Captain), Benjamin Jones (Treasurer), Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer), Conor McDermottroe (Public Relations Officer),

Absent(with apologies)

Absent Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 19 September 2001 1741 – 2000

Distribution DURC Committee

Overview of topics

1 Freshers Week

1 Freshers Week 1

2 Treasurer’s Report 1

3 A&E Report 2

4 Fees 2

5 Printing 2

6 RCMS 2

7 Open Day 3

8 Non-Cap and Cap grants 3

9 SGM 3

1 Freshers Week

John wanted to know if we needed to open the range in the mornings. Alan and Conor agreed that it was unnecessary as it was unlikely that people would be interested before noon. It was agreed that the range would only open between 1200 and 1600 each day. Task: Try to get PC working before Freshers Week (Mark)

Steven will do the talk for the introductory meeting and John will do the posing in full kit for the meeting.

2 Treasurer’s Report

£400 approx. in bank account. Bills are incoming and we have little or no ammunition. Alan is to pay his RO credits tomorrow. Grant meetings are coming up and non-capital budget decisions need to be made soon if we are to apply on the 11th.

3 A&E Report

We need to order ammunition. Discussion followed on the nature of this purchase - do we need Lapua Master for training or will T22 suffice? The consensus was that the difference in quality was smaller than anyone could take advantage of at present. Mark was skeptical of this, but agreed that T22 was the better buy.

The topic of bulk purchasing of ammunition was brought up. A fund is needed to buy bulk lots of ammunuition. Social spending cutbacks are necessary and were discussed. Internal competitions will involve BYOB receptions or whip-rounds instead of clubsponsored beverages, as this approach worked very well at the air rifle championships.

4 Fees

Buying ammunition from John Walsh was discussed, and John is to inquire more into this. EARC ammunuition logistics were discussed and John is again to inquire into this. Tracking procedures for ammunition were also reviewed. Task: Enquire as to ordering ammunition from JK Walsh (John)

It was decided to order 3000 rounds of Winchester T22 and 1000 of Lapua Master. Task: Order 3000 rounds T22, 1000 rounds Master (John)

4 Fees

The DUCAC levy is to rise to two euros this year. This prompts a discussion of the membership fee. Raising it to £3 was suggested, and reasonably well received, but it was decided to confer with other clubs before a final decision is made.

Ammunition prices for members were discussed. ROs will be expected to shoot T22 to lead by example.

5 Printing

We must start printing leaflets, membership forms and membership cards. It was agreed that the number on the membership card would be the receipt book number. Ben will tidy up the boxes and put in a hash sign to ensure clarity in the design of the cards, and send the design to Alan for printing.

Mark must also book the lecture theatre so we know where and when the introductory meeting is to be held so that can be put on the leaflets. Task: Book lecture theatre for Introductory Meeting (Mark)

6 RCMS

Mark is to send the RCMS tables to Conor so he can work on the website interface to the RCMS database. Task: Send RCMS tables to Conor (Mark)

7 Open Day

Ben reported that DUCACs original Open Week plan had been dropped in favour of an Open Day. Michelle Bennet is to be contacted. More details will be forthcoming from DUCAC later.

8 Non-Cap and Cap grants

Ben says we must decide on NonCapital grants budgeting before freshers’ week as the meeting is on Oct 11. It was decided to meet September 26th at 1800 on the range to

9 SGM

discuss NonCap budgeting and to discuss Capital budgeting in three months when the grant meeting is due.

9 SGM

An SGM is needed to adjust the constitiution as DUCAC are unhappy with some aspects of it. Mark asked if Keith’s working group had met yet. It had not. It was decided to handle the SGM after Freshers’ Week due to workload.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

9 SGM

Tasks agreed on at meeting

Mark Try to get PC working before Freshers Week (19 September 2001) [/136] /2 John Enquire as to ordering ammunition from JK Walsh (19 September

2001) [/169] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

John Order 3000 rounds T22, 1000 rounds Master (19 September 2001) [/174] /2 Mark Book lecture theatre for Introductory Meeting (19 September

2001) [/200] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Send RCMS tables to Conor (19 September 2001) [/206] . . . . . . . /2